the perimeter of a rectangle is 148 m. The length is 2 m more than twice the width. what is the dimensions
Answer by checkley77(12844)

P=2L+2W
L=2W+2
148=2(2W+2)+2W
148=4W+4+2W
148=6W+4
6W=148-4
6W=144
W=14/6
W=24 ANS. FOR THE WIDTH.
L=2*24+2=48=2=50 ANS. FOR THE LENGTH.
PROOF:
148=2*50+2*24
148=100+48
148=148